# Gauge fields in the presence of the electroweak bubble wall

### v2

@@ Abstract
The gauge field theory of the standard electroweak model in the presence of the electroweak bubble wall is investigated with an eye toward its applications to microscopic phenomena, which are supposed to have occurred during the phase transition in the early universe. The asymptotic fields are defined anew so that the effects of the position-dependent Higgs condensate are taken into account through the position-dependent $W$ and $Z$ boson masses. A novel method of massive gauge field quantization in the $R_{\xi}$-gauge with $\xi=1$ is proposed for the case of the position-dependent masses. Our procedure is based on the eigenfunction expansion method associated with second-order differential operators, i.e., a sort of generalized Fourier expansion.The commutation relations of creation and annihilation operators of various wave propagation modes are given in terms of the so-called spectral function. The decoupling of unphysical states from the physical S-matrix is also investigated along the line of Kugo-Ojima’s quartet mechanism on the basis of the BRST symmetry. It is pointed out that one of the quartet fields is not merely the unphysical scalar field but should be a linear combination of the unphysical scalar and the gauge fields. The physical and unphysical polarizations of the gauge field waves are unambiguously distinguished and this will help us evaluate the friction caused by the physical polarization states of $W$ and $Z$ boson waves on the bubble wall during the phase transition in the early universe.
@@ 1 Introduction
The purpose of the present paper is to scrutinize how and to what extent the presence of the electroweak bubble modifies the framework and notions in gauge theories, such as asymptotic fields, field quantization method,  [^1] See Refs. boedeker2 - ai3 for other approaches to field quantization under bubble wall backgrounds. creation and annihilation operators, classification of vector field polarizations, the decoupling of unphysical states from the physical S-matrix, and so forth. It should be emphasized that these are not totally academic problems but are strongly motivated by the recent intense studies on the bubble wall expansion dynamics in the early universe.
@@ 2 The electroweak bubble wall
The trivial constant solutions to (11) are $v=0$ and another one $v=v_{0}\neq 0$ satisfying $V^{\prime}(v_{0}^{\>2}/2)=0$, which correspond to symmetry-restored and symmetry-broken vacuum solutions, respectively. If, however, $v$ is not a constant but depends on one of the three space coordinates, say $z$, then the equation (11) for $v=v(z)$ becomes a non-trivial ordinary differential equation,
@@ 3.1 Asymptotic fields
For now, we have to put our definition of the asymptotic fields on a sound basis by specifying, first of all, a mathematical method of constructing the retarded and advanced Green’s functions. Technical tools for elaborating the Green’s functions are prepared in the course of solving (23) and of getting all the wave propagation modes for given mass function $M_{Z}(z)$. Therefore we investigate the solutions to (23) first in Section 3.2 and then will come back later in Section 3.4 to the retarded and advanced Green’s functions.
@@ 3.3 Scalar field quantization in the presence of the electroweak bubble wall
3.4 The advanced and retarded Green’s functions
@@ 4.1 The gauge fixing and auxiliary fields
The point of rewriting ${\cal L}_{\rm GF}$ as in (62) is that the gauge fields in (63), (64) and (65) are arranged so that they turn out to be $W^{\pm}$-, $Z$- and the electromagnetic ($A_{\mu}$) fields which are related to the original gauge fields $A_{\mu}^{a}$ $(a=1,2,3)$ and $B_{\mu}$ via
$W_{\mu}^{\pm}\equiv\frac{A_{\mu}^{1}\mp iA_{\mu}^{2}}{\sqrt{2}}\>,\>\>\>Z_{\mu}\equiv\frac{gA_{\mu}^{3}-g^{\>\prime}B_{\mu}}{\sqrt{g^{2}+g^{\>\prime\>2}}}\>,\>\>\>A_{\mu}\equiv\frac{g^{\>\prime}A_{\mu}^{3}+gB_{\mu}}{\sqrt{g^{2}+g^{\>\prime\>2}}}\>,$ (68)
@@ 4.2 The Faddeev-Popov ghost Lagrangian
We just follow the standard procedure to write down the Faddeev-Popov Lagrangian ${\cal L}_{FP}$ associated with the gauge fixing (59), that is,
@@
On the ground of argument in Refs.kugoojima1 - kugoojima5 as regard to the (anti-)hermiticity properties of ghost and anti-ghost fields, we take
@@ 4.3 The BRST transformation
$\delta\lambda\left({\cal L}_{\rm GF}+{\cal L}_{\rm FP}\right)$ $=$ $\delta_{B}\bigg{\{}\sum_{a=1}^{3}\overline{c}^{a}\bigg{(}F^{a}-\frac{1}{2}{\cal B}^{a}\bigg{)}+\overline{c}^{0}\bigg{(}F^{0}-\frac{1}{2}{\cal B}^{0}\bigg{)}\bigg{\}}$ (81)
$-\delta_{B}\bigg{\{}\partial^{\mu}\bigg{(}\sum_{a=1}^{3}\overline{c}^{a}A^{a}_{\mu}+\overline{c}^{0}B_{\mu}\bigg{)}\bigg{\}}\>,$
@@ 5.2.2 Quantization of the $Z^{\mu}$ ($\mu=0,1,2$)-fields
$\left[\alpha^{(a)}_{i}(\vec{p}_{\perp},\lambda),\;\alpha^{(b)\>{\dagger}}_{j}(\vec{p}_{\perp}^{\>\prime},\lambda^{\>\prime})\right]$ $=$ $-\eta^{ab}\>\delta^{2}(\vec{p}_{\perp}-\vec{p}_{\perp}^{\>\prime})\>\delta(\lambda-\lambda^{\>\prime})\>\frac{d\rho_{ij}(\lambda)}{d\lambda}$
@@ 5.2.3 Coupled equations of $Z^{3}(t,\vec{x}_{\perp},z)$ and $\chi^{3}(t,\vec{x}_{\perp},z)$
Here $\widehat{\rho}_{kl}(\lambda)$ is a 4 $\times$ 4 matrix and describes the spectrum of $\lambda$. The integration $d\widehat{\rho}_{kl}(\lambda)$ is, precisely speaking, the Stieltjes integral. We, however, assume simply that the $\lambda$-spectrum is continuous and therefore we regard (188) as the usual Riemann integration. We are thus led to the expansion of the pair of the fields in (150) in the following way,
@@ 5.3 The auxiliary field ${\cal B}^{Z}(t,\vec{x}_{\perp},z)$
which is a desirable relation. We also have to confirm that ${\cal B}^{Z}(t,\vec{x}_{\perp},z)$ is a dynamical variable independent of $\chi^{3}(t,\vec{x}_{\perp},z)$ and the commuting property
@@ 7 Physical Polarization States
$\alpha_{i}^{(1)\>{\dagger}}(\vec{p}_{\perp},\lambda)|0\rangle\>,\hskip 28.45274pt\alpha_{i}^{(2)\>{\dagger}}(\vec{p}_{\perp},\lambda)|0\rangle\>,\hskip 28.45274pt(i=1,2)$ (265)
are positive metric states and also satisfy (87) and are therefore physical. In addition to these, we have so far discussed the following states
@@
$\sum_{k=1}^{4}\mathfrak{h}_{i}^{(k)}\beta_{k}^{{\dagger}}(\vec{p}_{\perp},\lambda)|0\rangle\hskip 28.45274pt(i=1,2)\;,$ (268)
@@
which is an algebraic equation to determine the coefficients $\mathfrak{h}^{(k)}_{i}$. In this way we classify the four states $\beta^{{\dagger}}_{k}(\vec{p}_{\perp},\lambda)|0\rangle,\>\>(k=1,\cdots,4)$ into a pair of states given in (267) and another pair of states (268). The latter pair of states contributes to the physical S-matrix, but the former does not. To sum up, physical three pairs of polarization states consist of (265) and (268).
@@ 8 Summary
Now that the theoretical framework of the standard model gauge theory in the presence of the bubble wall has been put into a good shape, we are now able to launch with renewed interest into reanalyses of plasma interactions with the electroweak bubble wall. We have already witnessed in the last several decades a large variety of publications boedeker1 , boedeker2 - ai3 , dine1 - garcia1 on the coupled dynamics between the bubble wall and the plasma. The microscopic approach based on quantum field theories as given in the present work is appropriate for the ultra-relativistic regime of the wall velocity. It will be interesting to see how the pressure exerted on the wall by plasma particles depends on the Lorentz boost factor, when the effects of multi-soft vector boson emission in the transition radiation are summed up somehow to all orders. A careful look at the infrared region of the phase space is required, and the framework given in this paper will be the most reliable, since the position dependent mass is taken into account without any approximation. For the case of relatively low velocity of the bubble wall, hydrodynamical methods under the local thermal equilibrium condition are often used in literatures. Plasma particles in the distribution function in such approaches are endowed with definite fixed masses and with definite three-momenta. It could be necessary that particles in the distribution function should be labelled with $(\vec{p}_{\perp},\lambda)$ instead of three-momenta, as we did in the present paper, so that the momentum transfer from the plasma particles to the bubble wall should be within the reach in the theoretical framework. We hope we come to these issues in our future publications.
@@ Appendix B The BRST transformation
The BRST transformation rules of Heisenberg fields given in Section 4.3 are here rewritten by using physical variables, (66), (67), (68), (73) and (74). (The same formulas were previously listed up in aoki2 by using their conventions.) The gauge and scalar field transformation formulas are as follows,

### v4

@@ Abstract
The gauge field theory of the standard electroweak model in the presence of the electroweak bubble wall is investigated with a view to its applications to microscopic phenomena, which are believed to have occurred during the phase transition in the early universe. The asymptotic fields are defined anew so that the effects of the position-dependent Higgs condensate are taken into account through the position-dependent $W$ and $Z$ boson masses. A novel method of massive gauge field quantization in the $R_{\xi}$-gauge with $\xi=1$ is proposed for the case of the position-dependent masses. Our procedure is based on the eigenfunction expansion method associated with second-order differential operators, i.e., a sort of generalized Fourier expansion. The commutation relations of creation and annihilation operators of various wave propagation modes are given in terms of what is known as the spectral function. The decoupling of unphysical states from the physical S-matrix is also investigated along the line of Kugo-Ojima’s quartet mechanism on the basis of the BRST symmetry. It is pointed out that one of the quartet fields is not merely the unphysical scalar field but should be a linear combination of the unphysical scalar and the gauge fields. The physical and unphysical polarizations of the gauge field waves are unambiguously distinguished and this will help us evaluate the friction caused by the physical polarization states of $W$ and $Z$ boson waves on the bubble wall during the phase transition in the early universe.
@@ 1 Introduction
The purpose of the present paper is to scrutinize how and to what extent the presence of the electroweak bubble modifies the framework of and notions in gauge theories, such as asymptotic fields, field quantization method,  [^1] See Refs. boedeker2 - ai3 for other approaches to field quantization under bubble wall backgrounds. creation and annihilation operators, classification of vector field polarizations, the decoupling of unphysical states from the physical S-matrix, and so forth. It should be emphasized that these are not totally academic problems but are strongly motivated by the recent intense studies on the bubble wall expansion dynamics in the early universe.
@@ 2 The electroweak bubble wall
By discarding the terms linear in the Higgs field, the separation between background and quantum fluctuations is well defined. The trivial constant solutions to (11) are $v=0$ and another one $v=v_{0}\neq 0$ satisfying $V^{\prime}(v_{0}^{\>2}/2)=0$, which correspond to symmetry-restored and symmetry-broken vacuum solutions, respectively. If, however, $v$ is not a constant but depends on one of the three space coordinates, say $z$, then the equation (11) for $v=v(z)$ becomes a non-trivial ordinary differential equation,
@@ 3.1 Asymptotic fields
For now, we have to put our definition of the asymptotic fields on a sound basis by specifying, first of all, a mathematical method of constructing the retarded and advanced Green’s functions that appear in (LABEL:eq:retardedgreensfunction) and (LABEL:eq:advancedgreensfunction). Technical tools for elaborating the Green’s functions are prepared in the course of solving (23) and of getting all the wave propagation modes for given mass function $M_{Z}(z)$. Therefore we investigate the solutions to (23) first in Section 3.2 and then will come back later in Section 3.4 to the retarded and advanced Green’s functions.
A remark to be inserted here for our better understanding is that, in (21) and (22), we take only the temporal infinity ($t\to\pm\infty$) and the spatial coordinates (in particular the $z$-coordinate) are kept fixed. This is of course the usual procedure in the LSZ formalism to define the asymptotic fields. The overall profile of the Higgs condensate $v(z)$ is unaltered as depicted in Figure 1 and the symmetry-broken and symmetry-restored regions coexist when we solve the wave equation (23). For those who would like to compute the friction on the bubble wall in the plasma rest frame, the bubble wall is expanding very rapidly and the Higgs condensate is filling almost all corners of the space in the infinite future. Therefore it might be puzzling for them to think of wave propagation modes of (23) which are streching not only in symmetry-broken but also symmetry-restored regions in the $t\to+\infty$ limit. In the present paper we are just aiming at the LSZ-like formulation in the presence of the Higgs condensate and the bubble wall rest frame turns out to be the most convenient amenable to the LSZ-like formulation. At present, however, we do not have much to say about other Lorentz frames for the use of the LSZ-like formulation.
@@ 3.3 Scalar field quantization in the presence of the electroweak bubble wall
3.4 The retarded and advanced Green’s functions
@@ 4.1 The gauge fixing and auxiliary fields
The point of rewriting ${\cal L}_{\rm GF}$ as in (62) is that the gauge fields in (63), (64) and (65) are arranged so that they turn out to be $W^{\pm}$-, $Z$- and the electromagnetic ($A_{\mu}$)-fields which are related to the original gauge fields $A_{\mu}^{a}$ $(a=1,2,3)$ and $B_{\mu}$ via
$W_{\mu}^{\pm}\equiv\frac{A_{\mu}^{1}\mp iA_{\mu}^{2}}{\sqrt{2}}\>,\>\>\>Z_{\mu}\equiv\frac{gA_{\mu}^{3}-g^{\>\prime}B_{\mu}}{\sqrt{g^{2}+g^{\>\prime\>2}}}\>,\>\>\>A_{\mu}\equiv\frac{g^{\>\prime}A_{\mu}^{3}+gB_{\mu}}{\sqrt{g^{2}+g^{\>\prime\>2}}}\>.$ (68)
@@ 4.2 The Faddeev-Popov ghost Lagrangian
We just follow the standard procedure to write down the Faddeev-Popov Lagrangian ${\cal L}_{\rm FP}$ associated with the gauge fixing (59), that is,
@@
On the ground of argument in Refs.kugoojima1 - kugoojima5 in regard to the (anti-)hermiticity properties of ghost and anti-ghost fields, we take
@@ 4.3 The BRST transformation
$\delta\lambda\left({\cal L}_{\rm GF}+{\cal L}_{\rm FP}\right)=\delta_{B}\Xi\>,$ (81)
$\Xi\equiv\sum_{a=1}^{3}\overline{c}^{a}\bigg(F^{a}-\frac{1}{2}{\cal B}^{a}\bigg)+\overline{c}^{0}\bigg(F^{0}-\frac{1}{2}{\cal B}^{0}\bigg)-\partial^{\mu}\bigg(\sum_{a=1}^{3}\overline{c}^{a}A^{a}_{\mu}+\overline{c}^{0}B_{\mu}\bigg)\>,$
@@ 4.4 The conserved current and the conserved charge
In the previous subsection, we have seen that ${\cal L}_{\rm GF}+{\cal L}_{\rm FP}$ can be expressed as in (81), which is equivalently rewriten as
$\delta\lambda\left({\cal L}_{\rm GF}+{\cal L}_{\rm FP}\right)=i\left[\delta\lambda Q_{B},\Xi\right]\>.$ (88)
The subsidiary condition (87) indicates
$\langle{\rm phys}|\left({\cal L}_{\rm GF}+{\cal L}_{\rm FP}\right)|{\rm phys}^{\prime}\rangle=0\>,$ (89)
which implies that the effects of adding $\left({\cal L}_{\rm GF}+{\cal L}_{\rm FP}\right)$ to the original symmetric Lagrangian by hand become null in the physical subspace and physical quantities are independent of the choice of the gauge. In the present paper, we set $\xi=1$ in the $R_{\xi}$ gauge, but the relations such as (88) and (89) are both valid independently of the way of gauge fixing. In fact the proof of the gauge independence of the S-matrix usually starts with (88), but we have to be careful about the gauge dependence of renormalized quantities. In the present paper we do not enter the details of the renormalization procedure in the non-trivial $v(z)$ case, and we will discuss the gauge independence of physical quantities after scrutinizing the renormalization in our separate publications.
@@ 5.2.2 Quantization of the $Z^{\mu}$ ($\mu=0,1,2$)-fields
$\left[\alpha^{(a)}_{i}(\vec{p}_{\perp},\lambda),\;\alpha^{(b)\>{\dagger}}_{j}(\vec{p}_{\perp}^{\>\prime},\lambda^{\>\prime})\right]$ $=$ $-\eta^{ab}\>\delta^{2}(\vec{p}_{\perp}-\vec{p}_{\perp}^{\>\prime})\>\delta(\lambda-\lambda^{\>\prime})\>\frac{d\rho_{ij}(\lambda)}{d\lambda}\>,$
@@ 5.2.3 Coupled equations of $Z^{3}(t,\vec{x}_{\perp},z)$ and $\chi^{3}(t,\vec{x}_{\perp},z)$
Here $\widehat{\rho}_{kl}(\lambda)$ is a 4 $\times$ 4 matrix and describes the spectrum of $\lambda$. The integration $d\widehat{\rho}_{kl}(\lambda)$ is, precisely speaking, the Stieltjes integral. We, however, assume simply that the $\lambda$-spectrum is continuous or equivalently we assume the absence of bound state solutions of the potential $U(z)$ in (157). If there would be bound state solutions, they could affect the calculation of the friction on the bubble wall, but for now we leave it for future investigations. Under this assumption we regard (190) as the usual Riemann integration. We are thus led to the expansion of the pair of the fields in (152) in the following way,
@@ 5.3 The auxiliary field ${\cal B}^{Z}(t,\vec{x}_{\perp},z)$
which is a desirable relation. We also have to confirm that ${\cal B}^{Z}(t,\vec{x}_{\perp},z)$ is a variable independent of $\chi^{3}(t,\vec{x}_{\perp},z)$ and the commuting property
@@ 6.4 The Slavnov-Taylor identity and the BRST quartet mechanism
6.5 BRST cohomology classes
A supplementary remark to be added here is that the zero norm property (264) can also be seen by rewriting (263) in the following way
$|(\vec{p}_{\perp},\lambda,i),(\vec{p}_{\perp}^{\>\prime},\lambda^{\>\prime},i^{\>\prime})\rangle=-Q_{B}\>\left(\mathfrak{n}_{i}^{{\dagger}}(\vec{p}_{\perp},\lambda)\overline{\gamma}_{i^{\prime}}^{{\dagger}}(\vec{p}_{\perp}^{\>\prime},\lambda^{\prime})|0\rangle\right)\>.$ (267)
This can be confirmed by noting $Q_{B}|0\rangle=0$ and the following commutation and anti-commutation relations
$\left[Q_{B},\mathfrak{n}_{i}^{{\dagger}}(\vec{p}_{\perp},\lambda)\right]=-2i\lambda\gamma_{i}^{{\dagger}}(\vec{p}_{\perp},\lambda),\>\>\>\>\left\{Q_{B}\>\overline{\gamma}_{i^{\prime}}^{{\dagger}}(\vec{p}_{\perp}^{\>\prime},\lambda^{\prime})\right\}=-\mathfrak{b}_{i^{\prime}}^{{\dagger}}(\vec{p}_{\perp}^{\prime},\lambda^{\prime}),$ (268)
which can be deduced, respectively, from (261) and (260). A little more generally one can confirm the identity
$\thinspace\left\{\mathfrak{n}_{i}^{{\dagger}}(\vec{p}_{\perp},\lambda)\mathfrak{b}_{i^{\>\prime}}^{{\dagger}}(\vec{p}_{\perp}^{\>\prime},\lambda^{\>\prime})+2i\lambda\>\gamma_{i}^{{\dagger}}(\vec{p}_{\perp},\lambda)\overline{\gamma}_{i^{\>\prime}}^{{\dagger}}(\vec{p}^{\>\prime}_{\perp},\lambda^{\>\prime})\right\}|{\rm phys}\rangle$ (269)
$=$ $-Q_{B}\>\left(\mathfrak{n}_{i}^{{\dagger}}(\vec{p}_{\perp},\lambda)\overline{\gamma}_{i^{\prime}}^{{\dagger}}(\vec{p}_{\perp}^{\>\prime},\lambda^{\prime})|{\rm phys}\rangle\right)$
for all physical states $|{\rm phys}\rangle$ because of $Q_{B}|{\rm phys}\rangle=0$. These are zero norm states and are orthogonal to all of the physical states. In gauge theories in general, two physical states, $|{\rm phys}\rangle$ and $|{\rm phys}^{\>\prime}\rangle$, are said to belong to the same BRST cohomology class if the difference between these two states are in the form of $Q_{B}|\psi\rangle$, i.e.,
$|{\rm phys}\rangle=|{\rm phys}^{\>\prime}\rangle+Q_{B}|\psi\rangle\>.$ (270)
The states (269) expressed in terms of creation operators all belong to the set of $Q_{B}|\psi\rangle$ states.
@@ 7 Physical Polarization States
$\alpha_{i}^{(1)\>{\dagger}}(\vec{p}_{\perp},\lambda)|0\rangle\>,\hskip 28.45274pt\alpha_{i}^{(2)\>{\dagger}}(\vec{p}_{\perp},\lambda)|0\rangle\>,\hskip 28.45274pt(i=1,2)\>,$ (271)
are positive norm states and satisfy the physical state condition (87). (Note the relations $\left[Q_{B},\alpha_{i}^{(1)\>\dagger}(\vec{p}_{\perp},\lambda)\right]=0$ and $\left[Q_{B},\alpha_{i}^{(2)\>\dagger}(\vec{p}_{\perp},\lambda)\right]=0$.) In addition to (271), we have so far discussed the following states
@@
$\sum_{k=1}^{4}\mathfrak{h}_{i}^{(k)}\beta_{k}^{{\dagger}}(\vec{p}_{\perp},\lambda)|0\rangle\>,\hskip 28.45274pt(i=1,2)\;,$ (274)
@@
which is an algebraic equation to determine the coefficients $\mathfrak{h}^{(k)}_{i}$. Eq. (278) states that the two sets of states, (273) and (274), are orthogonal to each other. In this way we classify the four states $\beta^{{\dagger}}_{k}(\vec{p}_{\perp},\lambda)|0\rangle,\>\>(k=1,\cdots,4)$ into a pair of states given in (273) and another pair of states (274). The latter pair of states contributes to the physical S-matrix, but the former does not. To sum up, physical three pairs of polarization states consist of (271) and (274).
@@ 8 Summary
Now that the theoretical framework of the standard model gauge theory in the presence of the bubble wall has been put into a good shape, we are now able to launch with renewed interest into reanalyses of plasma interactions with the electroweak bubble wall that have been dealt with in a large variety of publications boedeker1 , boedeker2 - ai3 , dine1 - doomsday in the last several decades . The microscopic approach as given in the present paper is appropriate for studying momentum transfer in $z$-direction from plasma to the bubble wall in a systemmatic way. Note that each wave propagation mode of the gauge boson is characterized by $\vec{p}_{\perp}$ and $\lambda$ and there is no conservation law associated with $\lambda$. It will be interesting to see how the pressure exerted on the wall by plasma particles depends on the Lorentz boost factor, when the effects of multi-soft vector boson emission in the transition radiation are summed up effectively to all orders. A careful look at the infrared region of the phase space is required, and the framework given in this paper will be the most reliable, since the position dependent mass is taken into account without any approximation.
There exists another type of approach based on hydrodynamical methods under the local thermal equilibrium condition. This approach is supposed to be useful for the case of relatively low velocity of the bubble wall. Plasma particles in the distribution function in such approaches are endowed with definite fixed masses and with definite three-momenta, and the momentum non-conservation in the $z$-direction is introduced in a somewhat ad hoc manner. In order to improve such a situation, the field theoretical approach as developed in the present paper should be incorporated in the hydrodynamical approach. The quantum version of the Boltzmann equation should be based on the finite temperature Green’s functions, which must be expressed in terms of $\vec{p}_{\perp}$ and $\lambda$ without using such an ill-defined quantity as the $z$-component of momenta. Also various techniques, that have been developed in non-equilibrium statistical physics, must be applied extensively to Green’s functions. We hope to come to these issues in our future publications.
@@ Appendix B The BRST transformation
The BRST transformation rules of Heisenberg fields given in Section 4.3 are rewritten here by using physical variables, (66), (67), (68), (73) and (74). (The same formulas were previously listed up in aoki2 by using their conventions.) The gauge and scalar field transformation formulas are as follows,